1. You have been given a sample that is a 1/50 dilution of the original. You measure 3.4 mg glucose in 0.5 ml of that sample. What is the concentration in the original sample? If you have 400 ml of that sample, how many grams of glucose is in the total sample?
Answer #1

The measured concentration of glucose is 3.4 mg / 0.5 mL sample.

The original sample is diluted 50 times so we would multiply the amount of glucose by 50: 3.4 * 50 = 170 mg / 0.5 mL original sample, or 170 * 2 = 340 mg / mL glucose in the original sample.

Here the dilution is 1 / 50 and so the dilution factor is 50.

We have 400 mL of the original sample. So the amount of glucose in the original sample would be: 340 * 400 = 136000 mg glucose, or 136000 / 1000 = 136 g glucose.